# Decoding and Engineering the Phytobiome Communication for Smart Agriculture

### v1

@@ I Introduction
The world population is estimated to reach nearly $10$ billion by $2050$, and accordingly, the demand to increase agricultural production is rising. While the $70\%$ of the world’s water is consumed in agriculture, water scarcity is increasing every year due to climate change. In addition, agrochemicals such as pesticides and fertilizers are used to crop production efficiency. The excessive use of these agrochemicals is harmful to the environment and living things from humans to bees. Furthermore, their delivery efficiency is low, e.g., less than $0.1\%$ of applied pesticides reach their target pests in the USA [1]. These issues raise concerns about public health, biodiversity loss, and food security. Hence, sustainable, efficient, and eco-friendly methods are needed in agriculture.
@@
To improve agricultural production efficiency, smart agriculture applications have emerged thanks to various technologies such as the Internet of Things (IoT), robotics, and machine learning/artificial intelligence (ML/AI) [2]. Via these applications, plants and their environment are monitored, typically focusing on environmental parameters. In addition, smart agriculture applications include plant monitoring with cameras and optical sensors to detect plant stress such as pest attacks, infections, and nutrient deficiencies. These collected sensor data can be transferred to monitoring centers using available communication technologies such as WiFi, ZigBee, or cellular networks, and processed to support decisions on resource management, including irrigation and agrochemicals.
@@
The drawback of these applications is their delayed detection of plant stress. For instance, a bacterial infection is typically detected only after visual symptoms appear, which increases crop loss. As a solution, electrophysiological signals, which are used by plants to internally communicate information, emerged for early stress detection [3]. However, this type of monitoring treats the plant as a black box, lacking insight into internal mechanisms. Moreover, current smart agriculture applications do not consider the plant and its environment as an ecosystem but only measure and extract information from individual components.
@@
Nevertheless, a plant is not independent of its environment and organisms around it. To establish a holistic perspective, the phytobiome, which is defined as the area including the plant, its environment, and organisms within and around the plant, has been proposed [4]. However, the intricate communication among the organisms in the phytobiome is far from being fully understood in terms of modeling and utilizing it for agricultural applications. This communication is essential for understanding the relationships among the organisms and for identifying ways to exploit them effectively. Recent progress in biophysics, molecular communication (MC), and mathematical biology highlight that information gathering and processing are fundamental to an organism’s fitness, i.e., its ability to adapt and survive [5]. Hence, there is a need to model and design the phytobiome communication in a unified framework.
@@
As illustrated in Fig. 1, phytobiome members such as bacteria, fungi, nematodes, and insects interact with plants by exchanging molecules, i.e., MC. MC emerged as a promising field that blends communication engineering with biology, biophysics, and chemistry. It was proposed as a paradigm for the communication among nanomachines (NMs) to accomplish advanced tasks such as targeted drug delivery [6]. Moreover, the MC concept gave rise to the Internet of Bio-Nano Things (IoBNT), where nanonetworks consisting of NMs are remotely controlled through the Internet. Importantly, MC has also been utilized to understand natural phenomena such as infectious disease spread [7] and information propagation in plants [8].
@@ II Decoding the Phytobiome Communication
A phytobiome is a network wherein a plant has intricate connections with various organisms as illustrated in Fig. 1. This network includes viruses and organisms across various kingdoms, encompassing bacteria, archaea, protists, fungi, and animals [9]. Communication through molecular and electrical signals is imperative for the survival and reproductive success of phytobiome inhabitants. In this section, we first review the state-of-the-art of this communication in two categories: intra-kingdom and inter-kingdom. Subsequently, we explain how the communication engineering perspective can be applied to model phytobiome communication in various scales, including experimental validation of our MC model for electrophysiological signals.
@@ II-A1 Intra-kingdom Communication
Plants also communicate with other plants, employing both “wireless” and “wired” communication as shown in Fig. 1 [10]. The former mechanism is employed to warn nearby plants via airborne MC signals known as volatile organic compounds (VOCs). For example, a tomato plant attacked by herbivores secretes VOCs, which diffuse in the air and are received by other tomato plants to induce their defense mechanisms via JA. The “wired” communication occurs via fungi around the roots. Mycorrhiza, defined as the symbiotic relationship between fungi and plants, plays an important role in enhancing the plant’s capability to find water and nutrients in exchange for metabolites produced by photosynthesis. The mycorrhizal fungal networks are employed to communicate the defense mechanism via JA as well as other mechanisms such as kin recognition. Thanks to the underground “wired” fungal networks, a forest can be called a “wood wide web”.
@@ II-A2 Inter-kingdom Communication
Inter-kingdom communication involves interactions of organisms across different kingdoms. Primarily, this communication is achieved by enhancing, mimicking, degrading, and inhibiting the intra-kingdom communication molecules, such as QS molecules, by organisms from other kingdoms. The QS mechanism plays a significant role in the symbiotic relationship of plants and bacteria. For example, plants can “hack” this bacterial communication by emitting rosmarinic acid, a QS mimicker molecule that binds to autoinducer receptors of bacteria [11]. Thus, plants induce an immature QS response and remove bacteria and their associated biofilm. In addition, VOCs emitted by bacteria and fungi can be sensed by plants to induce defense mechanisms, while plants can modulate fungal growth and mycotoxin production through secreted lipids.
Although microorganisms such as bacteria and fungi are generally perceived as harmful to plants, they provide advantages to plants through disease resistance, stress tolerance, and nutrient acquisition [4]. For instance, nitrogen-fixing bacteria convert atmospheric nitrogen into a usable form for plants. Plants can recruit microorganisms to help them resist stress, e.g., drought due to climate change. This occurs through root exudates facilitating MC between plants and microorganisms.
The defense mechanism of plants can also be induced by pheromones secreted by nematodes, causing infections. These pheromones can also be sensed by fungi to prey on nematodes. In addition, plants can affect the sexual and aggregation behaviors of insects by enhancing or inhibiting insect pheromones. Furthermore, the communication between the plant and viruses, archaea, and protists is not fully recognized, although there is evidence showing that these organisms may affect the function of other microbes and the plant defense mechanisms [4]. Moreover, the communication can be among more than two kingdoms, e.g., MC between the plant and the bacteria secreted by the Colorado potato beetle suppresses the plant defense via JA. Next, we elaborate on a framework to clarify this complex phytobiome communication.
@@ II-B Modeling the Phytobiome as a Communication Network
It is well understood that organisms use information and communication to increase their evolutionary fitness. Moreover, the fitness value of information can be quantified with information-theoretic analysis, such as Kelly betting, which assesses how information contributes to survival and adaptation. For instance, the maximum growth rate of an organism such as a plant in an uncertain environment depends on the entropy of the environmental state [5]. Communication in phytobiomes occurs across multiple scales, each with distinct dynamics. At the microscale, intra-organismal communication involves molecules like plant hormones transmitted through phloem tissues in response to stress. At the mesoscale, interkingdom interactions occur between plants and microbial communities via mechanisms such as bacterial QS and fungal communication at the root–soil interface. At the macroscale, communication spans across phytobiomes, e.g., stressed plants emit VOCs alerting neighboring plants. Analyzing phytobiomes as communication networks across these scales enables a deeper understanding of plant behavior and opens opportunities to design targeted interventions for enhanced agricultural efficiency, as detailed in Section III. This multi-scale perspective not only enables tractable decomposition of complex interactions but also aligns with the nested, modular organization of biological systems. To formalize this analysis, we propose a multi-scale communication framework illustrated in Fig. 2.
@@ II-B1 Microscale Communication
As depicted in Fig. 2, transmitter cells release molecules via passive diffusion or active transport mechanisms, such as ion pumps/channels, or exocytosis in extracellular vesicles. In phytobiomes, the communication channels differ from conventional MC models due to the unique biological structures of plants. For instance, plasmodesmata directly connect adjacent plant cells, enabling targeted intercellular transport, while vascular tissues like xylem and phloem support long-range advection-based communication. Furthermore, receiver cells detect these molecules via plasmodesmata, receptor binding, and endocytosis. While the core dynamics can be modeled with the advection-diffusion-reaction framework [12], several plant-specific challenges must be considered.
@@ II-B2 Mesoscale Communication
We define mesoscale communication as inter-kingdom communication among phytobiome members. As illustrated in Fig. 2, the plant is the center of the phytobiome communication, similar to a hub in a local area network (LAN). While communication paths among the plant, bacteria, fungi, and animal kingdoms are well-documented (solid lines in Fig. 2), paths involving archaea, protists, and viruses need further experimental studies (dashed lines in Fig. 2). Such communication paths ultimately affect plant health. For example, our study in [11] shows that disrupting the MC between bacteria and plants can enhance plant health by disrupting the bacterial biofilm. This MC perspective supports holistic system modeling of the phytobiome. Additionally, the analogy between communication networks and mesoscale phytobiome communication enables tools from network theory and science. For instance, connections in the phytobiome resemble a partial mesh network where the plant is the access point. Hence, tools from the graph theory can be used to reveal information propagation within the phytobiome.
@@ II-C1 Relation of MC to Action Potential Signals
Molecular exchange among plant cells generates electrophysiological signals through voltage changes across cell membranes. The flux of ions, such as Ca2+, through ion channels/pumps, as illustrated in Fig. 2, causes depolarization and repolarization of the cell membrane. These voltage changes occur in slow (mm/s) and fast (cm/s) fashions, referred to as variation potential and action potential (AP) signals, respectively. We focus on APs due to their faster propagation speed. AP signals are induced by stimuli or stress. For example, in Venus flytrap, APs trigger rapid leaf closure to trap insects. Similar AP signals occur in fungi, indicating their importance for potential applications, such as fungal computers [13].
MC enables AP generation among cells. Hence, the MC perspective can be utilized to model these electrophysiological signals. For instance, the voxel model proposed by our group in [8], where each cell is represented as a 3-D box and MC occurs through diffusion and active transport among cells, is employed to model APs. The application of communication and information theory allows us to quantify how reliably information is transmitted between cells via the mutual information between the transmitted signal and the receiver’s output signal. Additionally, the propagation speed is determined by measuring the time difference at which this mutual information surpasses a defined threshold. These insights help assess the efficiency of communication mechanisms in plants. Readers are referred to our work [8] for mathematical details.
Figure 3: Experimental setup for measuring AP signals in a Mimosa pudica plant. 1) Mimosa pudica plant, 2) Electrophysiological sensor, 3) Signal amplifier, 4) Single board computer. Adapted from [8].
@@ II-C2 Experimental Validation
To validate the MC-based voxel model of AP generation, we conducted experiments on Mimosa pudica, also known as the sensitive plant closing its leaves when touched. Both the model and its experimental validation were introduced in our work [8]. As illustrated in Fig. 3, the experimental setup consists of an electrophysiological sensor connected to the plant’s stem and soil, a signal amplifier, and a single-board computer to measure voltage changes in APs triggered by touch stimuli, which start the communication from the leaves to other plant organs, e.g., stem or root. Notably, the measurement of electrophysiological signals is prone to environmental interference, particularly from the electricity grid. To mitigate this source of noise, we applied notch filtering around $50$ Hz.
For multiple AP signals induced by two stimuli on the plant, the numerical results of the MC-based voxel model and measured signals, using the parameter values in [8], are given in Fig. 4. These results of our proof-of-concept study validate the ability of MC to model electrophysiological signals and pave the way towards understanding the phytobiome communication. The next section elaborates on novel applications facilitated by engineering the phytobiome communication.
Figure 4: Experimental and numerical results for AP signals, from [8].
@@ III-A Phytobiome Monitoring
The first application leverages electrophysiological signals originating from intercellular communication within the phytobiome (see Section II-C). Under stress conditions, e.g., drought, pest infestation, or nutrient deficiency, plants generate time-varying electrophysiological signals enabled by MC to coordinate responses such as the closure of stomata (tiny pores in the leaf’s skin used for gas exchange) or defense activation. These stress-induced signals exhibit characteristic changes in amplitude, frequency, and temporal dynamics, which are already being analyzed by ML/AI techniques in both academic and industrial settings for early stress detection. Fig. 5 (steps 1-4) illustrates this process, where electrophysiological signals are recorded via the sensor in the bio-cyber interface and then processed to diagnose the stress type. For example, in [3], features extracted from recorded time-series signals were classified using the Extreme Gradient Boosting ML method, achieving an accuracy of $80\%$. Deep learning methods, such as multi-layer perceptrons and convolutional neural networks, were also applied in the literature. These diagnostic results can be shared with remote farmers through the Internet for smart agriculture applications. However, how mesoscale phytobiome communication alters these signals is still unclear and requires further research to improve classification accuracy using signal processing and ML methods.
@@ III-B Targeted Delivery of Agrochemicals and Genes
This application integrates phytobiome monitoring with delivery in two modes: fully autonomous and semi-autonomous. In the fully autonomous mode (steps 1, 2, 5, and 6 in Fig. 5), the injection of NMs is decided autonomously via ML/AI based on diagnostic results for plant stress types, e.g., nutrient deficiency, infection, or pest infestation. In the semi-autonomous mode (steps 1-7 in Fig. 5), the injection decision is made by farmers through an internet connection. Afterward, NMs with their payloads localize their target cells by using MC among themselves and bind to them via engineered ligands such as proteins and aptamers. These target cells can be plant cells hosting herbivores or pathogenic bacteria. NMs can be configured to track plant stress hormones like JA to locate affected regions. For instance, pesticides can be delivered to leaf cells where herbivores are located, or NMs can be designed to target receptors on the pests’ surfaces for direct delivery. Fertilizers can also be similarly directed to nutrient-deficient cells through NMs. In both modes, the closed loop starting and ending in the bio-cyber interface enables continuous monitoring, allowing ML/AI methods to be fed with more data to increase the accuracy of stress diagnosis and injection decisions. Furthermore, gene transfer using modified DNA can enhance plant resilience. Nanoparticles can be employed to mediate DNA from NMs to target cells [1].

### v2

@@ I Introduction
The world population is estimated to reach nearly $10$ billion by $2050$, and accordingly, the demand to increase agricultural production is rising. While the $70\%$ of the world’s water is consumed in agriculture, water scarcity is increasing every year due to climate change. In addition, agrochemicals such as pesticides and fertilizers are used to improve crop production efficiency. The excessive use of these agrochemicals is harmful to the environment and living things from humans to bees. Furthermore, their delivery efficiency is low, e.g., less than $0.1\%$ of applied pesticides reach their target pests in the USA [vega2020nanoscale]. These issues raise concerns about public health, biodiversity loss, and food security. Hence, sustainable, efficient, and eco-friendly methods are needed in agriculture.
@@
To improve agricultural production efficiency, smart agriculture applications have emerged thanks to various technologies such as the Internet of Things (IoT), robotics, and machine learning/artificial intelligence (ML/AI) [shaikh2022recent]. Via these applications, plants and their environment are monitored, typically focusing on environmental parameters. In addition, smart agriculture applications include plant monitoring with cameras and optical sensors to detect plant stress such as pest attacks, infections, and nutrient deficiencies. These collected sensor data can be transferred to monitoring centers using available communication technologies such as WiFi, ZigBee, or cellular networks, and processed to support decisions on resource management, including irrigation and agrochemicals.
@@
The drawback of these applications is their delayed detection of plant stress. For instance, a bacterial infection is typically detected only after visual symptoms appear, which increases crop loss. As a solution, electrophysiological signals, which are used by plants to internally communicate information, emerged for early stress detection [najdenovska2021identifying]. However, this type of monitoring treats the plant as a black box, lacking insight into internal mechanisms. Moreover, current smart agriculture applications do not consider the plant and its environment as an ecosystem but only measure and extract information from individual components.
@@
Nevertheless, a plant is not independent of its environment and organisms around it. To establish a holistic perspective, the phytobiome, which is defined as the area including the plant, its environment, and organisms within and around the plant, has been proposed [trivedi2020plant]. However, the intricate communication among the organisms in the phytobiome is far from being fully understood in terms of modeling and utilizing it for agricultural applications. This communication is essential for understanding the relationships among the organisms and for identifying ways to exploit them effectively. Recent progress in biophysics, molecular communication (MC), and mathematical biology highlight that information gathering and processing are fundamental to an organism’s fitness, i.e., its ability to adapt and survive [donaldson2010fitness]. Hence, there is a need to model and design the phytobiome communication in a unified framework.
@@
As illustrated in Fig. 1, phytobiome members such as bacteria, fungi, nematodes, and insects interact with plants by exchanging molecules, i.e., MC. MC emerged as a promising field that blends communication engineering with biology, biophysics, and chemistry. It was proposed as a paradigm for the communication among nanomachines (NMs) to accomplish advanced tasks such as targeted drug delivery [farsad2016comprehensive]. Moreover, the MC concept gave rise to the Internet of Bio-Nano Things (IoBNT), where nanonetworks consisting of NMs are remotely controlled through the Internet. Importantly, MC has also been utilized to understand natural phenomena such as infectious disease spread [gulec2022mobile] and information propagation in plants [awan2019communication].
@@ II Decoding the Phytobiome Communication
A phytobiome is a network wherein a plant has intricate connections with various organisms as illustrated in Fig. 1. This network includes viruses and organisms across various kingdoms, encompassing bacteria, archaea, protists, fungi, and animals [leach2017communication]. Communication through molecular and electrical signals is imperative for the survival and reproductive success of phytobiome inhabitants. In this section, we first review the state-of-the-art of this communication in two categories: intra-kingdom and inter-kingdom. Subsequently, we explain how the communication engineering perspective can be applied to model phytobiome communication in various scales, including experimental validation of our MC model for electrophysiological signals.
@@ II-A1 Intra-kingdom Communication
Plants also communicate with other plants, employing both “wireless” and “wired” communication as shown in Fig. 1 [sharifi2021social]. The former mechanism is employed to warn nearby plants via airborne MC signals known as volatile organic compounds (VOCs). For example, a tomato plant attacked by herbivores secretes VOCs, which diffuse in the air and are received by other tomato plants to induce their defense mechanisms via JA. The “wired” communication occurs via fungi around the roots. Mycorrhiza, defined as the symbiotic relationship between fungi and plants, plays an important role in enhancing the plant’s capability to find water and nutrients in exchange for metabolites produced by photosynthesis. The mycorrhizal fungal networks are employed to communicate the defense mechanism via JA as well as other mechanisms such as kin recognition. Thanks to the underground “wired” fungal networks, a forest can be called a “wood wide web”.
@@ II-A2 Inter-kingdom Communication
Inter-kingdom communication involves interactions of organisms across different kingdoms. Primarily, this communication is achieved by enhancing, mimicking, degrading, and inhibiting the intra-kingdom communication molecules, such as QS molecules, by organisms from other kingdoms. The QS mechanism plays a significant role in the symbiotic relationship of plants and bacteria. For example, plants can “hack” this bacterial communication by emitting rosmarinic acid, a QS mimicker molecule that binds to autoinducer receptors of bacteria [gulec2023stochasticd]. Thus, plants induce an immature QS response and remove bacteria and their associated biofilm. In addition, VOCs emitted by bacteria and fungi can be sensed by plants to induce defense mechanisms, while plants can modulate fungal growth and mycotoxin production through secreted lipids.
Although microorganisms such as bacteria and fungi are generally perceived as harmful to plants, they provide advantages to plants through disease resistance, stress tolerance, and nutrient acquisition [trivedi2020plant]. For instance, nitrogen-fixing bacteria convert atmospheric nitrogen into a usable form for plants. Plants can recruit microorganisms to help them resist stress, e.g., drought due to climate change. This occurs through root exudates facilitating MC between plants and microorganisms.
The defense mechanism of plants can also be induced by pheromones secreted by nematodes, causing infections. These pheromones can also be sensed by fungi to prey on nematodes. In addition, plants can affect the sexual and aggregation behaviors of insects by enhancing or inhibiting insect pheromones. Furthermore, the communication between the plant and viruses, archaea, and protists is not fully recognized, although there is evidence showing that these organisms may affect the function of other microbes and the plant defense mechanisms [trivedi2020plant]. Moreover, the communication can be among more than two kingdoms, e.g., MC between the plant and the bacteria secreted by the Colorado potato beetle suppresses the plant defense via JA. Next, we elaborate on a framework to clarify this complex phytobiome communication.
@@ II-B Modeling the Phytobiome as a Communication Network
It is well understood that organisms use information and communication to increase their evolutionary fitness. Moreover, the fitness value of information can be quantified with information-theoretic analysis, such as Kelly betting, which assesses how information contributes to survival and adaptation. For instance, the maximum growth rate of an organism such as a plant in an uncertain environment depends on the entropy of the environmental state [donaldson2010fitness]. Communication in phytobiomes occurs across multiple scales, each with distinct dynamics. At the microscale, intra-organismal communication involves molecules like plant hormones transmitted through phloem tissues in response to stress. At the mesoscale, interkingdom interactions occur between plants and microbial communities via mechanisms such as bacterial QS and fungal communication at the root–soil interface. At the macroscale, communication spans across phytobiomes, e.g., stressed plants emit VOCs alerting neighboring plants. Analyzing phytobiomes as communication networks across these scales enables a deeper understanding of plant behavior and opens opportunities to design targeted interventions for enhanced agricultural efficiency, as detailed in Section III. This multi-scale perspective not only enables tractable decomposition of complex interactions but also aligns with the nested, modular organization of biological systems. To formalize this analysis, we propose a multi-scale communication framework illustrated in Fig. 2.
@@ II-B1 Microscale Communication
As depicted in Fig. 2, transmitter cells release molecules via passive diffusion or active transport mechanisms, such as ion pumps/channels, or exocytosis in extracellular vesicles. In phytobiomes, the communication channels differ from conventional MC models due to the unique biological structures of plants. For instance, plasmodesmata directly connect adjacent plant cells, enabling targeted intercellular transport, while vascular tissues like xylem and phloem support long-range advection-based communication. Furthermore, receiver cells detect these molecules via plasmodesmata, receptor binding, and endocytosis. While the core dynamics can be modeled with the advection-diffusion-reaction framework [jamali2019channel], several plant-specific challenges must be considered.
@@ II-B2 Mesoscale Communication
We define mesoscale communication as inter-kingdom communication among phytobiome members. As illustrated in Fig. 2, the plant is the center of the phytobiome communication, similar to a hub in a local area network (LAN). While communication paths among the plant, bacteria, fungi, and animal kingdoms are well-documented (solid lines in Fig. 2), paths involving archaea, protists, and viruses need further experimental studies (dashed lines in Fig. 2). Such communication paths ultimately affect plant health. For example, our study in [gulec2023stochasticd] shows that disrupting the MC between bacteria and plants can enhance plant health by disrupting the bacterial biofilm. This MC perspective supports holistic system modeling of the phytobiome. Additionally, the analogy between communication networks and mesoscale phytobiome communication enables tools from network theory and science. For instance, connections in the phytobiome resemble a partial mesh network where the plant is the access point. Hence, tools from the graph theory can be used to reveal information propagation within the phytobiome.
@@ II-C1 Relation of MC to Action Potential Signals
Molecular exchange among plant cells generates electrophysiological signals through voltage changes across cell membranes. The flux of ions, such as Ca2+, through ion channels/pumps, as illustrated in Fig. 2, causes depolarization and repolarization of the cell membrane. These voltage changes occur in slow (mm/s) and fast (cm/s) fashions, referred to as variation potential and action potential (AP) signals, respectively. We focus on APs due to their faster propagation speed. AP signals are induced by stimuli or stress. For example, in Venus flytrap, APs trigger rapid leaf closure to trap insects. Similar AP signals occur in fungi, indicating their importance for potential applications, such as fungal computers [adamatzky2018towards].
MC enables AP generation among cells. Hence, the MC perspective can be utilized to model these electrophysiological signals. For instance, the voxel model proposed by our group in [awan2019communication], where each cell is represented as a 3-D box and MC occurs through diffusion and active transport among cells, is employed to model APs. The application of communication and information theory allows us to quantify how reliably information is transmitted between cells via the mutual information between the transmitted signal and the receiver’s output signal. Additionally, the propagation speed is determined by measuring the time difference at which this mutual information surpasses a defined threshold. These insights help assess the efficiency of communication mechanisms in plants. Readers are referred to our work [awan2019communication] for mathematical details.
Figure 3: Experimental setup for measuring AP signals in a Mimosa pudica plant. 1) Mimosa pudica plant, 2) Electrophysiological sensor, 3) Signal amplifier, 4) Single board computer. Adapted from [awan2019communication].
@@ II-C2 Experimental Validation
To validate the MC-based voxel model of AP generation, we conducted experiments on Mimosa pudica, also known as the sensitive plant closing its leaves when touched. Both the model and its experimental validation were introduced in our work [awan2019communication]. As illustrated in Fig. 3, the experimental setup consists of an electrophysiological sensor connected to the plant’s stem and soil, a signal amplifier, and a single-board computer to measure voltage changes in APs triggered by touch stimuli, which start the communication from the leaves to other plant organs, e.g., stem or root. Notably, the measurement of electrophysiological signals is prone to environmental interference, particularly from the electricity grid. To mitigate this source of noise, we applied notch filtering around $50$ Hz.
For multiple AP signals induced by two stimuli on the plant, the numerical results of the MC-based voxel model and measured signals, using the parameter values in [awan2019communication], are given in Fig. 4. These results of our proof-of-concept study validate the ability of MC to model electrophysiological signals and pave the way towards understanding the phytobiome communication. The next section elaborates on novel applications facilitated by engineering the phytobiome communication.
Figure 4: Experimental and numerical results for AP signals, from [awan2019communication].
@@ III-A Phytobiome Monitoring
The first application leverages electrophysiological signals originating from intercellular communication within the phytobiome (see Section II-C). Under stress conditions, e.g., drought, pest infestation, or nutrient deficiency, plants generate time-varying electrophysiological signals enabled by MC to coordinate responses such as the closure of stomata (tiny pores in the leaf’s skin used for gas exchange) or defense activation. These stress-induced signals exhibit characteristic changes in amplitude, frequency, and temporal dynamics, which are already being analyzed by ML/AI techniques in both academic and industrial settings for early stress detection. Fig. 5 (steps 1-4) illustrates this process, where electrophysiological signals are recorded via the sensor in the bio-cyber interface and then processed to diagnose the stress type. For example, in [najdenovska2021identifying], features extracted from recorded time-series signals were classified using the Extreme Gradient Boosting ML method, achieving an accuracy of $80\%$. Deep learning methods, such as multi-layer perceptrons and convolutional neural networks, were also applied in the literature. These diagnostic results can be shared with remote farmers through the Internet for smart agriculture applications. However, how mesoscale phytobiome communication alters these signals is still unclear and requires further research to improve classification accuracy using signal processing and ML methods.
@@ III-B Targeted Delivery of Agrochemicals and Genes
This application integrates phytobiome monitoring with delivery in two modes: fully autonomous and semi-autonomous. In the fully autonomous mode (steps 1, 2, 5, and 6 in Fig. 5), the injection of NMs is decided autonomously via ML/AI based on diagnostic results for plant stress types, e.g., nutrient deficiency, infection, or pest infestation. In the semi-autonomous mode (steps 1-7 in Fig. 5), the injection decision is made by farmers through an internet connection. Afterward, NMs with their payloads localize their target cells by using MC among themselves and bind to them via engineered ligands such as proteins and aptamers. These target cells can be plant cells hosting herbivores or pathogenic bacteria. NMs can be configured to track plant stress hormones like JA to locate affected regions. For instance, pesticides can be delivered to leaf cells where herbivores are located, or NMs can be designed to target receptors on the pests’ surfaces for direct delivery. Fertilizers can also be similarly directed to nutrient-deficient cells through NMs. In both modes, the closed loop starting and ending in the bio-cyber interface enables continuous monitoring, allowing ML/AI methods to be fed with more data to increase the accuracy of stress diagnosis and injection decisions. Furthermore, gene transfer using modified DNA can enhance plant resilience. Nanoparticles can be employed to mediate DNA from NMs to target cells [vega2020nanoscale].